OpenStudy (anonymous):

find the centre and radius of the circle with equation given by x^2+3x+y^2-2y-3=0

OpenStudy (anonymous):

Just Keep the general equation ax^2 + by^2 +2gx+2fy+c=0

OpenStudy (anonymous):

Centre = (-g,-f)

OpenStudy (anonymous):

x^2+3x+y^2-2y-3=0 C( -3/2 , 1)

OpenStudy (anonymous):

x^2 + 3x + (9/4) + y^2 - 2y + 1 = 3 + (9/4) + 1 [x+(3/2)]^2 + (y-1)^2 = 25/4 We get centre = (-3/2,1) Radius = 5/2

OpenStudy (anonymous):

radius = sqrt ( g^2 + f^2-C)